OpenStudy (anonymous):

Calculate the amount of simple interest on $9,000 deposited at an interest rate of 5% for two year A) $900 B) $922.50 C) $9900 D) 9922.50

OpenStudy (paki):

simple... use formula... I=P*R*T/100 I= 9000*5*2/100 = 900$... so option A.... @noah4151

OpenStudy (anonymous):

ok thanks @paki i wasnt there for class when we learned it and its do tommorow

OpenStudy (paki):

ahhh... welcome dear... :-)

OpenStudy (anonymous):

im a guy... @paki

OpenStudy (paki):

so what... i have no interest with gender... i am here to help everyone dear... :-) @noah4151

OpenStudy (anonymous):

just the word dear is usally used for girls but ok @paki

OpenStudy (paki):

ahhh... but in my country, this word is used for "respect purpose" @noah4151